# Improving Data Efficiency for LLM Reinforcement Fine-tuning Through Difficulty-targeted Online Data Selection and Rollout Replay

### v1

@@ D.1 Design and Implementation Details
The adapter is a GELU-activated MLP with three hidden layers, each containing 896 units and a dropout rate of 0.1. A LayerNorm is applied to the projection output to stabilize training. The calibration head is a two-layer MLP that takes the mean and standard deviation of reference set difficulties as input. The first output passes through a Softplus activation to yield the scale parameter $w^{(t)}$, while the second is transformed by a Tanh activation to produce a bounded bias term $b^{(t)}$, as defined in Section 4.1.
We collect training data from a set of LLMs that are disjoint from our policy models. These include Qwen2.5-Instruct and Qwen2.5-Math-Instruct series [46], Eurus-2-7B-PRIME [5], Mathstral-7B-v0.1444https://huggingface.co/mistralai/Mathstral-7B-v0.1, DeepSeek-R1-Distill-Qwen-1.5B [9], DeepScaleR-1.5B-Preview [27], and Qwen2.5-7B-SimpleRL-Zoo [50]. For each model, we sample query questions and reference questions from math datasets and compute their adaptive difficulty as supervision labels. Each training instance consists of a query question $q$, a reference set $\{(q_{i},d_{i})\}_{i=1}^{K}$ with known difficulty scores, and a ground-truth difficulty label $d_{q}$. Repeating this procedure across models yields the training dataset $\mathcal{D}_{\text{pred-train}}$.
@@
$\mathcal{L}_{\text{BCE}}=-\frac{1}{|\mathcal{D}_{\text{pred-train}}|}\sum_{(q,\{(q_{i},d_{i})\},d_{q})\in\mathcal{D}_{\text{pred-train}}}\left[d_{q}\log\hat{d}_{q,\mathrm{cal}}+(1-d_{q})\log(1-\hat{d}_{q,\mathrm{cal}})\right],$
@@
D.2 Qualitative Examples
Tab. 3 presents a qualitative example from the Qwen2.5-3B model, showing one unlabeled question alongside reference questions with the highest and lowest attention scores. The example demonstrates that our difficulty prediction framework assigns higher attention to reference questions that share key mathematical topics and structures (e.g., rhombus, incircle), while down-weighting unrelated questions.
Table 3: Qualitative example illustrating similarity-based attention mechanism in adaptive difficulty prediction. The table shows one unlabeled question and its top- and bottom-ranked reference questions by attention score. High-attention references (red) typically share similar concepts and difficulty with the target question (e.g., rhombus and incircle geometry), while low-attention references (blue) diverge in topic and are significantly easier.
@@ D.2 Qualitative Examples
Unlabeled Question [adaptive diff. $=1.000$, predicted score $=0.907$]
@@
# Attention Score Adaptive Diff. Reference Question
@@
Appendix E Implementation Details
E.1 Training Datasets and Models
Our experiments involve three model sizes: Qwen2.5-Math-1.5B, Qwen2.5-3B, and Qwen2.5-Math-7B [46]. We adopt four open-source datasets of mathematical reasoning for RL fine-tuning:
@@ E.1 Training Datasets and Models
  • DeepMath-103K [12]: A large-scale dataset focused on high-difficulty mathematical problems, constructed with rigorous data decontamination procedures to support reliable benchmark evaluation. We also sample 8,192 problems for training.
E.2 RL Fine-tuning Details
@@
Number of Steps 60
@@
E.3 Implementation Details of DOTS and RR
@@
Table 6: Hyperparameters of DOTS and RR.
@@
E.4 Evaluation Details
Consistent with RL fine-tuning, we use a sampling temperature of 0.6, top-p of 0.95, and the same prompt template. We evaluate model performance on four commonly-used mathematical reasoning benchmarks and report the average accuracy to mitigate benchmark-specific variance. We exclude benchmarks with very few questions, such as AIME 24 (30 questions) and AMC 23 (40 questions), as their limited size leads to high evaluation variance and unreliable performance comparisons for smaller models [14].
@@ E.4 Evaluation Details
Appendix F Additional Experimental Results
F.1 Ablation Study on the Adaptive Difficulty Prediction Framework
@@
We evaluate a baseline that directly uses frozen embeddings from the Qwen2.5-Math-1.5B-Instruct model without any training or calibration. In contrast, our framework incorporates trained adapter layers and a calibration head. As shown in Tab. 7, our framework consistently achieves significantly higher Pearson correlation with ground-truth adaptive difficulty across all settings. The poor performance of the off-the-shelf baseline highlights the necessity of further adapter layers and calibration for accurately predicting question difficulty.
@@
Model Dataset Off-the-shelf Embedding Our Framework (With Adapter Layers + Calibration)
@@
We further investigate the impact of the reference set size $K$ in RL fine-tuning. Fig. 6 compares the performance of the original GRPO and the DOTS method under reference set sizes of 128 and 256, trained with Qwen2.5-Math-1.5B and Qwen2.5-3B on the DeepScaleR dataset. The results show that a reference set size of 128 yields RL performance comparable to that of 256. This indicates that our approach is robust to smaller reference sets, enabling more efficient rollout collection without sacrificing RL fine-tuning quality.
Figure 7: Comparison between DOTS (ours) and an external difficulty-based curriculum baseline. The curves show average performance aggregated over four benchmarks with exponential smoothing for visualization. Note that the x-axis is the number of steps (rather than time). Our method consistently outperforms the baseline.
F.2 Case Study: Online Data Selection Via External Difficulty-based Curriculum
We additionally implement an online data selection baseline that relies on external difficulty annotations. Specifically, we use the DeepScaleR dataset and label each question with GPT-4o-mini, following the difficulty annotation prompt introduced in Luo et al. [27]. Each question is annotated 32 times, and the average score is used as its final difficulty.
We then follow a staged curriculum: in the first third of training steps, batches are sampled from the easiest third of the dataset; in the middle third, from the medium-difficulty third; and in the final third, from the hardest third. To ensure a fair comparison of online data selection strategies, we compare this baseline with DOTS (without RR). As shown in Fig. 7, our DOTS method consistently outperforms this baseline on both Qwen2.5-Math-1.5B and Qwen2.5-3B. We further discuss in the main text the limitations of such approaches, including the high cost of annotation and limited adaptability due to their reliance on fixed, hand-crafted curricula.

### v2

@@ C.1 Design and Implementation Details
The adapter is a GELU-activated MLP with three hidden layers, each containing 896 units and a dropout rate of 0.1. A LayerNorm is applied to the projection output to stabilize training. The calibration head is a two-layer MLP that takes the mean and standard deviation of reference set difficulties as input. The first output passes through a Softplus activation to yield the scale parameter $w^{(t)}$, while the second is transformed by a Tanh activation to produce a bounded bias term $b^{(t)}$.
We collect training data from a set of LLMs that are disjoint from our policy models. These include Qwen2.5-Instruct and Qwen2.5-Math-Instruct series [46], Eurus-2-7B-PRIME [5], Mathstral-7B-v0.1555https://huggingface.co/mistralai/Mathstral-7B-v0.1, DeepSeek-R1-Distill-Qwen-1.5B [9], DeepScaleR-1.5B-Preview [27], and Qwen2.5-7B-SimpleRL-Zoo [50]. For each model, we sample query questions and reference questions from math datasets and compute their adaptive difficulty as supervision labels. Specifically, each training instance consists of a query question $q$, a reference set with known difficulty scores $\{(q_{i},d_{i})\}_{i=1}^{K}$, and a ground-truth difficulty label $d_{q}$. Repeating this procedure across models yields the training dataset $\mathcal{D}_{\text{pred-train}}$.
@@
$\mathcal{L}_{\text{BCE}}=-\frac{1}{|\mathcal{D}_{\text{pred-train}}|}\sum_{(q,\{(q_{i},d_{i})\}_{i=1}^{K},d_{q})\in\mathcal{D}_{\text{pred-train}}}\left[d_{q}\log\hat{d}_{q,\mathrm{cal}}+(1-d_{q})\log(1-\hat{d}_{q,\mathrm{cal}})\right],$
@@
C.2 Qualitative Examples
Tab. 3 presents a qualitative example from the DeepScaler dataset using Qwen2.5-3B as the policy model, showing one unlabeled question along with the reference questions receiving the highest and lowest attention scores. The example demonstrates that our difficulty prediction framework assigns higher attention to reference questions that share key mathematical topics and structures (e.g., rhombus, incircle), while down-weighting unrelated questions.
Table 3: Qualitative example illustrating the similarity-based attention mechanism in adaptive difficulty prediction. The table shows one unlabeled question along with its top- and bottom-ranked reference questions by attention score. High-attention references (red) typically share similar concepts and difficulty with the target question (e.g., rhombus and incircle geometry), while low-attention references (blue) diverge in topic and are substantially easier.
@@ C.2 Qualitative Examples
Unlabeled Question [ground truth adaptive difficulty =1.000=1.000, predicted difficulty =0.907=0.907]
@@
# Attention Score Adaptive Difficulty Reference Question
@@
Appendix D Implementation Details
D.1 Training Datasets and Models
Our experiments involve three model sizes: Qwen2.5-Math-1.5B, Qwen2.5-3B, and Qwen2.5-Math-7B [46]. We adopt four open-source mathematical reasoning datasets for RL fine-tuning:
@@ D.1 Training Datasets and Models
  • DeepMath-103K [12]: A large-scale dataset focused on high-difficulty mathematical problems, constructed with rigorous data decontamination procedures to support reliable benchmark evaluation. We sample 8,192 problems for training.
  • SCP-25K [liu2025prorl, lu2025scp]: A large-scale dataset focused on STEM problems. We sample 8,192 problems for training.
D.2 RL Fine-tuning Details
@@
Number of Training Steps 60
Number of Gradient Steps 480
@@
D.3 Implementation Details of DOTS and RR
@@
Table 6: Hyperparameters of DOTS and $R$.
@@
D.4 Evaluation Details
Consistent with RL fine-tuning, we use a sampling temperature of 0.6, top-p of 0.95, and the same prompt template. We evaluate model performance on four commonly-used mathematical reasoning benchmarks and report the average accuracy to mitigate benchmark-specific variance.
@@ D.4 Evaluation Details
  • MMLU Science: We sample all problems in the field of physics, chemistry and biology from MMLU benchmark.
We exclude benchmarks with very few questions, such as AIME 24 (30 questions) and AMC 23 (40 questions), as their limited size leads to high evaluation variance and unreliable performance comparisons for smaller models [14]. We further justify this exclusion by evaluating the original GRPO on AIME 24 across various LLM-dataset combinations. Specifically, each of the 30 AIME 24 questions is evaluated 8 times, and the average accuracy (avg@8) is computed at regular intervals during training. As shown in Table 7, the accuracy fluctuates considerably across training steps without a clear upward trend. This high variance across steps underscores the difficulty of obtaining reliable evaluation signals on such small-scale datasets, especially for smaller models with limited reasoning capacity.
Table 7: Accuracy of original GRPO on AIME 24 across training steps. Each of the 30 questions is evaluated 8 times, and avg@8 accuracy is reported every 10 training steps. The results show high variance without clear trends, which limits evaluation reliability especially for smaller models.
Steps Qwen2.5-Math-1.5B + DeepScaler Qwen2.5-3B + DeepScaler Qwen2.5-Math-7B + DeepScaler
10 10.42 7.50 20.42
20 15.83 6.25 24.17
30 15.83 8.75 24.17
40 11.67 9.17 23.33
50 16.67 7.50 25.42
60 13.75 5.42 20.00
Appendix E Additional Experimental Results
E.1 Ablation Study on the Adaptive Difficulty Prediction Framework
@@
We evaluate a baseline that directly uses frozen embeddings from the Qwen2.5-Math-1.5B-Instruct model without any further training or calibration. In contrast, our framework incorporates trained adapter layers and a calibration head. As shown in Tab. 8, our framework consistently achieves significantly higher Pearson correlation with the ground-truth adaptive difficulty across all settings. The poor performance of the off-the-shelf baseline highlights the necessity of further adapter layers and calibration for accurately predicting question difficulty.
@@
Model Dataset Off-the-shelf Embedding Our Method (With Adapter Layers + Calibration)
@@
We further investigate the impact of the reference set size $K$ in RL fine-tuning. Fig. 7 compares the performance of the original GRPO and DOTS under reference set sizes of 128 and 256, using Qwen2.5-Math-1.5B and Qwen2.5-3B on the DeepScaleR dataset. The results show that a reference set size of 128 yields RL performance comparable to that of 256. This indicates that DOTS is robust to smaller reference sets, enabling more efficient rollout collection without sacrificing RL fine-tuning quality.
E.2 Additional Experiment with Extended Training Horizon
To further verify the stability of our findings, we extend training to 100 training steps (600 gradient steps) under two settings: Qwen2.5-Math-1.5B + DeepScaleR and Qwen2.5-3B + DeepMath. Notably, our method continues to outperform the original GRPO baseline.
Figure 8: Extended training to 100 training steps for two settings: Qwen2.5-Math-1.5B + DeepScaleR and Qwen2.5-3B + DeepMath. Our method consistently outperforms the original GRPO baseline.
E.3 Additional Results under Different Evaluation Views
In the main text, Fig. 3 presents performance over wall-clock time, while Fig. 5 uses training steps. For completeness, we provide alternate versions: Fig. 9 shows the step-based view corresponding to Fig. 3, and Fig. 10 shows the time-based view for Fig. 5.
Figure 9: Performance curves with training steps as x-axis. Replot of Fig. 3 using training steps.
Figure 10: Performance curves with wall-clock time as x-axis. Replot of Fig. 5 using wall-clock time.
Across both views—training steps and wall-clock time—DOTS+RR and DOTS consistently demonstrate strong performance, confirming the robustness of our improvements regardless of presentation format. Interestingly, as shown in Fig. 10, although DOTS (without RR) incurs a small overhead from reference rollouts and difficulty prediction, it requires substantially fewer training steps to reach the same final accuracy as the original GRPO. As a result, the overall training time is often reduced despite the per-step overhead.